Ted is not particularly creative. He uses the pickup line, "If I could rearrange the alphabet, I would put U and I together." The random variable x is the number of girls Ted approaches before encountering one who reacts positively.

XP(X)

1 0.01

2 0.2

3 0.4

4 0.3

What is the mean of the random vailable X?

What is the standard deviaition of the random variable X?
Answer #1

There is a typo in this question. P(X = 1) should be equal to 0.1 not 0.01. Because if that value is 0.01 then total sum of all probability is not going to be equal to one.

a) Mean = E(X) = 1 * 0.1 + 2 * 0.2 + 3 * 0.4 + 4 * 0.3 = 2.9

b) E(X2) = 12 * 0.1 + 22 * 0.2 + 32 * 0.4 + 42 * 0.3 = 9.3

Var(X) = E(X2) - (E(X))2 = 9.3 - 2.92 = 0.89

Stnadard deviation = sqrt(0.89) = 0.9438